OpenStudy (anonymous):

matrix equation help? {{12,7},{5,3}}x={{2,-1},{3,2}}

OpenStudy (amistre64):

an augment might be beneficial

OpenStudy (amistre64):

unless you know how to work an inverse? Ax = b A'Ax = A'b Ix = A'b x = A'b

OpenStudy (anonymous):

how can i plug in matrices?

OpenStudy (amistre64):

what are you working with? ti83?

OpenStudy (anonymous):

No. I meant how can I plug in matrices on Openstudy

OpenStudy (anonymous):

So I can show the equation correctly

OpenStudy (amistre64):

its labor intensive, but i did make an app for it, if you are using a laptop or desktop that is

OpenStudy (amistre64):

\[\left[ \begin{array}c 12 & 7\\5 & 3\\\end{array} \right] x= \left[ \begin{array}c 2 & -1\\3 & 2\\\end{array} \right]\]

OpenStudy (anonymous):

\[\[\left[ \begin{array}c 12 & 7\\5 & 3\\\end{array} \right]\] X = \[\left[ \begin{array}c 2 & -1\\3 & 2\\\end{array} \right]\]\]

OpenStudy (anonymous):

hows that?

OpenStudy (anonymous):

yeah

OpenStudy (amistre64):

it was designed to post one matrix hence the oddity of trying multiples, but knowing the latex sorts it out

OpenStudy (amistre64):

so, how do we approach this? inverse are augment? either way is fine. inverse may be simpler in this case tho

OpenStudy (amistre64):

we can augment it as this, and row reduce. \[\left[ \begin{array}c 12 & 7 & | & 2 & -1\\5 & 3 & | & 3 & 2\\\end{array} \right]\]

OpenStudy (anonymous):

how do you reduce it

OpenStudy (amistre64):

elementary row operations for course, get the left side worked into the identity matrix. and the right side works itself into x

OpenStudy (amistre64):

divide row1 by 12 and multiply by -5 and add to row2 12,7,2,-1 1,7/12,2/12,-1/12 -5,-35/12,-10/12,5/12 5, 3, 3, 2 -5,-35/12,-10/12, 5/12 5, 36/12, 36/12,24/12 ---------------------- 0, 1/12, 26/12, 29/12 multiply by 12 0, 1, 26, 29 new setup 1, 7/12, 2/12, -1/12 0, 1 , 26 , 29 multiply row 2 by -7/12 and add it to row1 1, 7/12, 2/12, -1/12 0, -7/12 , -26(7)/12 , -29(7)/12 -------------------------------- 1, 0 , -90/6 , -204/12 1, 0 , -15 , -17 1, 0, -15, -17 0, 1, 26, 29 I x

OpenStudy (amistre64):

did you try for an inverse?

OpenStudy (amistre64):

-35 12 7 5 3 36 det = 1 by some basic shortcut, A' has a pattern swap corners and negate the other 2 A' = 3 -7 -5 12